40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Mon Oct 14 17:01:43 2019, anchal, Notes, Laser, South Laser Switched Off! SouthLaserStopped.pdf
    Reply  Mon Oct 14 18:30:14 2019, anchal, Notes, PMC, South PMC Error Signal Oscillations SPMCMixerOutOscillations.pdf
       Reply  Tue Oct 15 18:54:16 2019, anchal, Notes, PMC, South PMC further problems 
          Reply  Wed Oct 16 16:56:43 2019, anchal, Notes, PMC, South PMC won't work with it's own Servo Card only 
             Reply  Wed Oct 16 17:33:08 2019, Koji, Notes, PMC, South PMC won't work with it's own Servo Card only 
                Reply  Wed Oct 16 18:46:58 2019, anchal, Notes, PMC, South PMC started working somehow :( 
                   Reply  Wed Oct 23 17:39:44 2019, anchal, Notes, PMC, South PMC started working somehow :( 
                      Reply  Fri Oct 25 16:50:12 2019, anchal, Notes, PMC, South PMC problem debugging efforts 
                         Reply  Tue Oct 29 19:10:59 2019, anchal, Notes, PMC, South PMC problem debugging efforts SPMC_Diagnostics.pdf
                            Reply  Wed Oct 30 15:15:28 2019, anchal, Notes, PMC, Replacing U5A AD602 didn't solve the issue. 
                               Reply  Wed Oct 30 18:35:31 2019, anchal, Notes, PMC, Issue closed 
                                  Reply  Mon Nov 4 12:15:01 2019, anchal, Notes, PMC, Changing the autolocking method 
Message ID: 2458     Entry time: Wed Oct 16 18:46:58 2019     In reply to: 2457     Reply to this: 2464
Author: anchal 
Type: Notes 
Category: PMC 
Subject: South PMC started working somehow :( 

Yeah, the cables are isolated and no inversion could happen.

Quote:

Are you sure that all the cables involved are isolated and there is no polarity inversion? e.g. The unfunctional combination provides HV to GND directry at the cable, for example.

 

Even more bizzarething is that it works now!  indecisionI'm not halucinating here but the same thing was not working before.I even have lab notes from yesterday when it wasn't working.

This is pretty bad as I don't want to be unaware of something in the lab that caused this. The only other clue here in all this is that the laser intensity changed. We control the intensity of light going into the experiment at (24,110) through the half waveplate before a PBS. Rana told me that the polarization direction of laser coming from NPRO shouldn't change, but since the incident of last Friday, I have had to change the rotation of this half waveplate in both directions to ensure same amound of light reaches the cavity as on North. since no alignment was changed at the south PMC, there is no reason for the mode matching to change drastically there or during the day. but this is the only fishy clue I've got for now.


Presently

Both cavities are locked. with same amount of gains in the FSS and PMC loops in both paths as before. The can's temperature as reached to the setpoint and the beatnote frequency PID is working to take it to 27.34 MHz. I'll set trigger for tonight for beatnote frequency noise measurement if the frequency reaches in the range of the photodiode.We then will know what is the impact on the experiment noise.

ELOG V3.1.3-